OpenStudy (anonymous):

Given a circle with a radius of 4, which is the circumference?

OpenStudy (anonymous):

\[C = 2 π r \]

OpenStudy (anonymous):

c = 2* 3.14r

OpenStudy (anonymous):

C=2*3.14r which would make the answer 25.2 if you round.

OpenStudy (anonymous):

14. C = 2 pi (4) Circumference = 8 pi feet 15. Same equation except given diameter except of radius. 10 pi yards 16. Circumference =Diameter pi 62.8/3.14 = 20 meters 17. Area = pi r^2 pi 5^2 = 25 pi